OpenStudy (anonymous):

Evaluate the polynomial for x= -1 -4x^3 + 9x^2 - 9x-2

OpenStudy (anonymous):

16

jhonyy9 (jhonyy9):

-4(-1) +9(-1)*2 -9(-1) -2 = 4+9+9-2=20

OpenStudy (anonymous):

Thank you

jhonyy9 (jhonyy9):

please if you like it always

OpenStudy (anonymous):

what?

jhonyy9 (jhonyy9):

the help